OpenStudy (anonymous):

the derivative of x^(-2x) is -2x^(-2x) (log(x)+1) why is there that +1?

OpenStudy (mathteacher1729):

See image. 1) Go to http://www.wolframalpha.com 2) type "derivative x^(-2x) Hope this helps. :)

OpenStudy (anonymous):

\[x^{-2x}\] take the log get \[-2x\ln(x)\] take the derivative using product rule get \[-2x\frac{1}{x} -2ln(x)\]

OpenStudy (anonymous):

now factor out the 2 get \[2(\ln(x)-1)\]

OpenStudy (anonymous):

of course the x's in the first term cancel. then multiply by the original function to get your answer: \[x^{-2x}(2(ln(x)-1)=2\times x^{-2x}(ln(x)-1)\]

OpenStudy (anonymous):

i made a mistake is see. start with \[-2-2ln(x)\] gives \[-2(ln(x)+1)\]

OpenStudy (anonymous):

sorry about that.

OpenStudy (anonymous):

but that is where the -2 comes from

OpenStudy (anonymous):

you can do this because ln(x) '=1/x so if you take the log and at the end you multiply by the function than you get the derrivate

OpenStudy (anonymous):

this is a question :)

OpenStudy (anonymous):

oh yes. if you want to take the derivative of a function with the variable in the exponent the easiest thing to do is first take the log.

OpenStudy (anonymous):

I have never seen this before, but it is a clever idea

OpenStudy (anonymous):

you know that the derivative of \[ln(f(x))\] is \[\frac{f'(x)}{f(x)}\]

OpenStudy (anonymous):

yep, I see

OpenStudy (anonymous):

which means you take the log, take the derivative, and then multiply by original function.

OpenStudy (anonymous):

yes yes, cool technique

OpenStudy (anonymous):

you can also view it this way: \[x^{2x}=e^{-2x\ln(x)}\]

OpenStudy (anonymous):

then take the derivative using the chain rule. the derivative of e to the something is just e to the something times the derivative of something. you will see that the work is identical

OpenStudy (anonymous):

thanks, I see why this helps

OpenStudy (anonymous):

in any case all the work is taking the derivative of the log of that thing, either way. hope this helps